40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Thu Jun 28 23:19:49 2018, johannes, Disha, Notes, Cavity, Gaussian beam mode for the new cavities New_Cavities.pdf
    Reply  Sun Jul 1 23:19:25 2018, johannes, Notes, Cavity, Gaussian beam mode for the new cavities 6x
       Reply  Tue Jul 3 11:58:25 2018, johannes, Notes, Cavity, Gaussian beam mode for the new cavities cavity_support_ring_gen2_15deg.PDFcavity_support_shelf.PDFshelf__contacts.pngcyl__contacts.png
          Reply  Wed Jul 11 22:38:54 2018, johannes, Notes, Cavity, Some assembly required 20180706_212353.jpg20180706_202052.jpg
             Reply  Wed Jul 11 23:27:38 2018, johannes, Notes, Cavity, new breadboard assembly 20180710_143617.jpg20180710_143629.jpg20180710_145112.jpg20180710_163221.jpg20180710_163803.jpg
                Reply  Wed Jul 11 23:51:18 2018, johannes, Notes, Cavity, first light 
                   Reply  Thu Jul 19 23:19:01 2018, johannes, Summary, Cavity, cryostat closed and pumping 20180719_114735.jpg20180717_164839.jpg
                      Reply  Sun Jul 22 00:52:10 2018, johannes, Summary, Cavity, transmission paths set up, mode-matching improved, post-baking ringdowns 6x
Message ID: 2105     Entry time: Sun Jul 22 00:52:10 2018     In reply to: 2104
Author: johannes 
Type: Summary 
Category: Cavity 
Subject: transmission paths set up, mode-matching improved, post-baking ringdowns 

Attachment #1 shows a picture of the modified pump line arrangement. The turbo unit was detached from the main pump body and bolted to a breadboard that is attached to the surrounding frame, hovering over the transmission setup. There are no more signs of overheating and I turned off the extra fan. After observing for a little while longer it was still not warming up like before, so I assume the load on the pump was simply too large in the beginning and pumping alternatingly the inner and outer volumes separately may have helped with its initial struggle. The extra fan will be removed. After a little over 48 hours of pumping we're closing in on 30 microtorr (#3), which is similar to pre-cavity-swap pumpdowns.

I aligned the transmitted beams onto DC PDs, CCD Camera, and Beat PD. Characterization will continue, writing down some initial observations:

  1. The natural beat frequency is around 470 MHz, as expected from the in-air case (#4)
  2. A cold start (lasers off -> lasers on + locked) does not cause a strong drift of cavity frequency anymore. With the old cavities it would take less than a minute for the laser control signal to rail, now it's good for 10+ minutes.
  3. The relative drift of the cavities is also much reduced, even in the current non-thermally controlled state.
  4. I don't see any strange polarization behavior (at least at room temperature). The light level on the gold PDs changes by only miniscule amounts when rotating the polarization hitting the cavity with a HWP after the FI. Transmitted polarization looks pretty pure as well.

Attachment #2 shows a schematic of the complete setup. The most significant change is that we're using the 1st order deflected beam of the AOM which completely extinguishes when the drive signal is cut. I aligned the AOM, maximizing the deflected power while bringing the passing beam as close as possible to the PZT transducer. I used this to make the first post-contacting and post-baking ringdown measurements of the new cavities. I measured ~13 ppm loss in the west, and ~14 ppm in the east. During the assembly the lowest loss I reported was 9 ppm, however I noticed I was using the length of the spacer (36.5 mm) to calculate the round-trip time, not the 37 mm if the mirror polish is included. Thid corrects the 9 pmm estimate to 11 ppm. I'm drrinh marginally increased loss in the cavities, which could have come from small mirror shifts during the contacting, and affixing in the tensioned baking jig, or the baking. So it's not still not bad at all but it means the caivities are slightly undercoupled.

I proceeded with working on the mode-matching and in the end observed dips of the reflected light of 96% relative depth. The under-coupling of the cavities reduces the observable dip by 1%, so the mode overlap is more like 97% for both paths.

Attachment 1: 20180720_204253.jpg  3.341 MB  Uploaded Sun Jul 22 02:04:06 2018  | Hide | Hide all
20180720_204253.jpg
Attachment 2: CryoTableDrawing3.pdf  118 kB  Uploaded Sun Jul 22 02:04:21 2018  | Hide | Hide all
CryoTableDrawing3.pdf
Attachment 3: pressure_trend.png  23 kB  Uploaded Sun Jul 22 02:31:26 2018  | Hide | Hide all
pressure_trend.png
Attachment 4: beat_trend.png  20 kB  Uploaded Sun Jul 22 02:31:50 2018  | Hide | Hide all
beat_trend.png
Attachment 5: ringdown_west.pdf  117 kB  Uploaded Sun Jul 22 02:51:48 2018  | Hide | Hide all
ringdown_west.pdf
Attachment 6: ringdown_east.pdf  116 kB  Uploaded Sun Jul 22 02:52:07 2018  | Hide | Hide all
ringdown_east.pdf
ELOG V3.1.3-